An electron moving in a magnetic field induces the magnetic B =

5.10-3T, in the direction that matches the unidirectional magnetic field line α = 60o

. Energy of

electron is equal to W = 1.64.10-16J. In this case


1
Expert's answer
2021-01-24T14:24:15-0500

... the magnetic force is


"F=qvB\sin60°"


"E=mv^2/2\to v=\sqrt{2E/m}="


"=\sqrt{2\cdot 1.64\cdot10^{-16}/9.1\cdot10^{-31}}=18.98\cdot10^6(m/s)"


"F=qvB\sin60°="


"=1.6\cdot10^{-19}\cdot18.98\cdot10^6\cdot 5\cdot10^{-3}\cdot \sin60°=1.3\cdot10^{-14}(N)"